a balloon ries from rest on the ground with constant accleration g/8 . a stone is dropped ffom a balloon when the ballioon has risen to a height of H m . the time taken by the stone to reach the ground is

Grade:Upto college level

1 Answers

AskiitianExpert Shine
14 years ago

Hi

Applying v2 -u2 = 2as ,, gives v2=2*g/8*H  , for upward motion of balloon. therefore velocity of stoned while dropped is v= (gh)0.5/2

Let the time taken by stone to reach ground is T. apply eq, s=ut + 1/2 a t2 whr s=H , u=-v , t=T , a=g

solve the quadratic equation to get the roots and take the positive root, hence answer is 2(h/g)0.5 .